Question 28 Error sum of squares (ESS) is computed as o a. È(9-Y,)? ob. Ŝ(v;-P) c. ][9:-x;) od. [9:-Y) O

Question 30 The estimated value of Y1 is given by o a. f = B, +BX ob. 9 = bo +byx oc. 9= B. + BAX +8 od. 8,= bo +bX+8Question 31 The adjusted R2 statistic a. is equal to the value of unadjusted R? o b.adjusts R2 for the degrees of freedom inR2 is calculated as o a. RSS/ESS ob.RSS/TSS oc. ESS/TSS od. 1 - (RSS/TSS)
Answer #1

28) The Total sum of squares (SST), Sum of squares due to regression (SSR) and Sum of squares due to errors (SSE) are given by:

2 SST ΣΥ, - Υ) SSE = Σ Σ Y, - Υ = i-1 SST i-1 SSE SSR

Hence, correct option is (b)

30) The estimated value with k independent variables is expressed as Y, = bo + b, X 11 + b2x + ... +62X 21 ki

For one independent variable the estimated value will be \hat {Y_i}=b_0+b_1X_1

Hence, correct option is (b)

31) The relation between adjusted R and R is given by

R 2 Adj = 1- (1 - R?)(n-1) (n - k-1), where k denotes degrees of freedom.

Adjusted R is always less than or equal to R .

Hence, correct option is (b)

Last question) R is calculated as (Regression sum of squares) / (Total sum of squares)

Hence, correct option is (b)